# A Survey on Medical Large Language Models: Technology, Application, Trustworthiness, and Future Directions

### v1

@@ Abstract
Large language models (LLMs), such as GPT series models, have received substantial attention due to their impressive capabilities for generating and understanding human-level language. More recently, LLMs have emerged as an innovative and powerful adjunct in the medical field, transforming traditional practices and heralding a new era of enhanced healthcare services. This survey provides a comprehensive overview of Medical Large Language Models (Med-LLMs), outlining their evolution from general to the medical-specific domain (i.e, Technology and Application), as well as their transformative impact on healthcare (e.g., Trustworthiness and Safety). Concretely, starting from the fundamental history and technology of LLMs, we first delve into the progressive adaptation and refinements of general LLM models in the medical domain, especially emphasizing the advanced algorithms that boost the LLMs’ performance in handling complicated medical environments, including clinical reasoning, knowledge graph, retrieval-augmented generation, human alignment, and multi-modal learning. Secondly, we explore the extensive applications of Med-LLMs across domains such as clinical decision support, report generation, and medical education, illustrating their potential to streamline healthcare services and augment patient outcomes. Finally, recognizing the imperative and responsible innovation, we discuss the challenges of ensuring fairness, accountability, privacy, and robustness in Med-LLMs applications, where ethical considerations, rigorous evaluation methodologies, and the formulation of regulatory frameworks are pivotal to fostering trustworthiness in these systems. Finally, we conduct a concise discussion for anticipating possible future trajectories of Med-LLMs, identifying avenues for the prudent expansion of Med-LLMs, such as technological progress and policy evolution. By consolidating above-mentioned insights, this review seeks to provide a comprehensive investigation of the potential strengths and limitations of Med-LLMs for professionals and researchers, ensuring a responsible landscape in the healthcare setting.
@@
Medical Large Language Model, Clinical Application, Trustworthy and Ethical Considerations.
@@ I Introduction
The emergence of foundation models has sparked a transformative wave within the Artificial Intelligence (AI) community in recent years. The success, particularly their remarkable generalization capacity across a myriad of downstream tasks, is attributed to significantly large model sizes and large-scale pre-training on expansive datasets [1, 2]. Such progression has witnessed the extraordinary evolution of large language models (LLMs) within the Natural Language Processing (NLP) society, characterized by a series of seminal advancements, such as BERT [3], T5 [4], and the latest GPT-4 [5]. Leveraging these powerful models, AI systems can now excel in generating human-like responses, hinting at the potential for highly sophisticated and interactive applications.
@@
§ V, for tree= top_class [Applications, for tree=fill=yellow!45,apply [Medical Diagnosis, Clinical Report Generation, Medical Education, Medical Robotics, Medical Language Translation, apply_wide] ] [Challenges , for tree=fill=yellow!45,apply [Protected Health Information, Clinical Workflows, Safety and Accountability,apply_wide] ] ] [Trustworthy
@@
Figure 1: Architecture of this survey.
@@
Overall, through these advanced explorations in the field of Med-LLMs, various effective perspectives are delivered to promote the rapid development of medical AI societies. From the technical point, it is still a basic manner to improve the performance of Med-LLMs via increased model scale and complexity, which is consistent with the LLM’s scaling law [77]. The Med-LLMs, often pre-trained and fine-tuned on massive biomedical literature corpora like PubMed [78], outperform general-purpose language models on domain-specific tasks. To adapt to diverse clinical tasks, fine-tuning Med-LLMs over an array of specialized medical tasks is required. Essentially, to track the potential pathways for developing Med-LLMs, several reviews [79, 80] primarily focus on exploring the clinical capabilities and the practical applications of Med-LLMs, including Electronic Health Records (EHRs) [81], health education [82], and diagnostic decision-making [83].
@@
Figure 2: The development history of LLMs is a story of rapid progress in AI and NLP. The overall model evolution is marked by innovations in statistical NLP (e.g., data-driven models), neural networks (e.g., RNNs and LSTMs), and self-attention (e.g., transformer). Throughout this journey, the development of LLMs has been fueled by advances in hardware, algorithmic innovation, and the availability of large-scale datasets.
@@ II Background and Technology
II-A Development History
The development history of LLMs charts a course through the progress in AI and NLP, promoted by the growth of computational power (i.e., GPUs) and the availability of vast datasets. We derive a detailed timeline highlighting the key milestones and breakthroughs for such evolution.
II-A1 Foundations and Early NLP (Pre-2000s)
  • the 1950s: The Turing test, originated from the imitation game by Alan Turing in 1950, is an evaluation metric to determine whether a machine can exhibit human-like intelligence, laying the theoretical groundwork for AI [85].
  • the 1980s-1990s: Rule-based systems dominate NLP, such as expert systems and the development of part-of-speech tagging [86]. At this time, statistical NLP begins to emerge, marking a shift towards data-driven models.
  • the 1990s: Recurrent Neural Networks (RNNs) are introduced, capable of handling sequential data, but face difficulties with vanishing gradients for long sequences [87].
  • 1997: LSTM (Long Short-Term Memory) units are proposed to address RNN’s limitations, enabling better learning of long-term dependencies [88].
II-A2 The Deep Learning Revolution (2010s)
  • 2010s: The popularity of deep learning, inspired by GPUs and large datasets, leads to breakthroughs in computer vision and speech recognition.
  • 2013: Word2Vec [89], the concept of word embedding, allows words to be represented as dense vectors capturing semantic and syntactic properties.
  • 2017: Vaswani et al. [90] introduced the Transformer architecture, which replaces recurrence with self-attention mechanisms, enabling parallelization and scalability.
  • 2018: OpenAI’s GPT (Generative Pre-trained Transformer)  [91] showcases the power of unsupervised pre-training on large datasets, followed by task-specific fine-tuning.
II-A3 Explosion of Pre-Trained LLMs
  • 2019: BERT (Bidirectional Encoder Representations from Transformers)  [3] revolutionizes NLP by using masked language modeling for pre-training, achieving state-of-the-art performance across various tasks.
  • 2020: GPT-3  [92], developed by OpenAI, with large parameters, demonstrates the potential of extremely large-scale models for zero-shot and few-shot learning, sparking industry-wide interest and debates about the future of AI.
  • 2021-Present: The trend towards even larger models continues with the introduction of models like GPT-4 [93].
Throughout this journey, LLMs have transformed from rudimentary rule-based systems to sophisticated neural architectures capable of generating coherent text, understanding context, and even displaying elements of creativity. The focus now is not only on increasing model capacity but also on responsible deployment, interpretability, and sustainability.
II-B Basic Architecture
@@
II-B1 Encoder-Decoder Structure
@@
II-B2 Self-Attention Mechanism
@@ II-B3 Positional Encoding
II-B4 Residual Connections & Layer Normalization
For each layer, residual connections are employed to mitigate the vanishing gradient problem during backpropagation. After the residual connection, layer normalization is applied to stabilize and speed up the training process. This structure ensures that the model can effectively learn from very deep architectures without suffering from degradation issues.
II-C Basic Training Techniques
II-C1 Pre-training (PT)
@@
Next Word Prediction (NWP) [95]. NWP, also known as Language Modeling (LM), aims to predict the most possible next word in a sequence given the previous words. At its core, NWP relies on understanding the context provided by the preceding words. This context helps the model infer the statistical likelihood of different words following in the sequence, i.e., given a huge corpus, the entire LLM is optimized with maximum likelihood estimation (MLE).
Mask Language Modeling (MLM) [3]. MLM involves training a model to predict the original identities of certain words (tokens) in a sentence that have been randomly replaced with a special mask token (often “[MASK]”). By predicting the missing words based on remained context, the model can learn to understand the semantic and bidirectional relationships between words. The basic masking strategy involves masking 15% of the tokens in a sequence. Not all masked tokens are treated equally. For the masked tokens, 80% are replaced by the “[MASK]” token, 10% are replaced by a random word, and 10% are left unchanged. This variation prevents the model from simply memorizing the masked positions and forces it to learn more generalizable context representations.
@@
Replaced Token Detection (RTD) [97]. RTD is a pre-training methodology designed to enhance language models’ capability to discern subtle semantic variations in the text by detecting tokens that have been deliberately substituted with synonyms or semantically equivalent terms. This approach leverages the richness of linguistic synonyms and near-synonyms to create a learning environment that sharpens the model’s sensitivity to the nuances of meaning. During training, two types of sentence instances are generated: (1) Unaltered Sentences: A portion of sentences remain unmodified, serving as a baseline for the model to learn the standard usage of words in context. (2) Altered Sentences: The remaining sentences undergo a careful replacement of selected words with synonyms, creating alternative versions that maintain the overall message but differ in lexical choice. By integrating RTD into the pre-training regimen, models can acquire a more nuanced understanding of semantic similarity and distinction. This refined understanding proves beneficial for downstream tasks that demand high levels of semantic precision, such as text entailment, sentiment analysis, and semantic textual similarity tasks, where the ability to differentiate between subtle shifts in meaning is crucial.
@@
Next Sentence Prediction (NSP) [3]. Punctuation naturally segments text, offering a logical foundation for developing pre-training strategies based on this inherent structure. The NSP task is designed to train models to discern whether a pair of sentences presented to them are contiguous segments taken from a text corpus. In practice, during NSP’s pre-training phase, two kinds of pairs are constructed: (1) Continuous Pairs (IsNext): Approximately 50% of the time, the model is given sentence pairs that are genuinely consecutive in the original text. (2) Random Pairs (NotNext): For the remaining 50%, the second sentence in the pair is randomly selected from elsewhere in the corpus, forming unrelated or discontinuous sequences. This balanced approach ensures that models not only learn to recognize sequential sentence structures but also develop an understanding of contextual continuity versus disjointed pairings.
@@
Sentence Order Prediction (SOP) [98]. As a pre-training task, SOP can improve a model’s understanding of the coherence and narrative flow in a text by determining the correct sequential order of sentences, which is particularly useful in training models for tasks that require comprehension of long-form text, dialogue understanding, or any scenario where the sequence of events matters. The primary goal of SOP is to teach models to recognize the logical progression of ideas within a text by predicting whether a pair of sentences are arranged in their original, sequential order or not. Thus, two types of sentence pairs are formulated: (1) Sequential Pairs: Half of the time, sentence pairs are kept in their original order as they appear in the text. These form the positive examples, labeled as ‘IsOrder’. (2) Non-Sequential Pairs: The remaining half are constructed by swapping the order of sentences from different parts of the text, creating incoherent or unrelated sequences. These are labeled as ‘NotOrder’, serving as negative examples. The model is fed these sentence pairs and tasked with predicting whether the given order is correct according to the narrative flow. In this way, SOP encourages models to learn the underlying logic and coherence that connects sentences, enabling them to infer the correct sequence of events or thoughts.
II-C2 Fine-tuning
Supervised Fine-tuning (SFT) [99, 100]. SFT follows a structured approach to adapt pre-trained models to specific tasks by further training them on new, labeled datasets. In SFT, the pre-trained LLM is fine-tuned on labeled datasets using supervised learning techniques.
Instruction Fine-tuning (IFT) [101]. IFT, also known as prompt-based fine-tuning or instruction tuning, is a recent adaptation of the traditional fine-tuning method, particularly popularized in the context of LLMs. It focuses on teaching models to follow instructions or prompts to perform a variety of tasks without explicit task-specific architecture changes or separate training phases for each task. Instead of fine-tuning on a single-task dataset, instruction fine-tuning involves using a diverse set of instructions or prompts that cover multiple tasks within the same training process. Besides, a medical IFT variant is proposed called Instruction Prompt Tuning (IPT) to train Med-PaLM.
Parameter-Efficient Fine-ting (PEFT) [102]. PEFT refers to techniques designed to adapt large pre-trained models to specific downstream tasks while minimizing the introduction of new parameters or modifying only a fraction of the existing ones. This approach aims to retain most of the pre-training knowledge, reduce computational costs, and alleviate the risk of overfitting, especially when the target task dataset is small. One popular and classical PEFT strategy is LoRA [103], which introduces low-rank matrices that are multiplied with the weight matrices of specific layers during forward and backward passes. This allows the model to learn task-specific adjustments without changing the original weights directly, reducing the number of new parameters needed.
II-C3 Reinforcement Learning from Human Feedback (RLHF)
II-C4 In-Context Learning (ICL)
ICL [106], also known as few-shot learning for LLMs, refers to the capability of generalizing and adapting to new tasks by providing examples within the prompt, without requiring additional fine-tuning on a specific training dataset. This method leverages the model’s pre-existing knowledge and its ability to understand the context to infer what is being asked and generate appropriate responses.
@@ III-A NLP Tasks under Medical Domain
III-A1 Medical Information Extraction
Medical Information Extraction (Med-IE) is a fundamental field to extract relevant medical information from unstructured or semi-structured sources such as electronic health records (EHRs), clinical notes, medical research articles, and other healthcare documents. The aim is to transform this textual data into structured formats that can be easily analyzed, shared, and utilized in various applications like decision support systems, population health management, and medical research. The common applications of Med-IE are summarized in Table II.
@@ III-A1 Medical Information Extraction
III-A2 Medical Question Answer
The Medical Question Answering (Med-QA) task is a specific application for building AI systems (e.g., LLMs) capable of understanding and answering complex medical questions from patients, healthcare providers, or researchers. The system typically leverages large amounts of structured and unstructured data sources such as electronic health records, medical textbooks, research articles, and online resources to generate accurate and reliable answers. Here, we summary how the Med-QA system works in the Table II. Information retrieval could involve searching through medical literature, patient histories, or other knowledge bases, which aims to identify and extract key pieces of evidence that are pertinent to the question.
III-A3 Medical Natural Language Inference
Medical Natural Language Inference (Med-NLI) is to deal with understanding the relationship between two pieces of medical text and determining whether one sentence logically entails, contradicts, or is neutral with respect to another. This task is also known as Recognizing Textual Entailment (RTE) in the broader context of NLP. In Med-NLI, the inputs are typically pairs of sentences, where one is called the premise and the other is the hypothesis. For example:
@@ III-A3 Medical Natural Language Inference
III-A4 Medical Text Generation
Medical Text Generation (Med-Gen), specifically focusing on tasks like text summarization, is an application of artificial intelligence where algorithms process and condense lengthy medical documents into concise summaries. This is particularly useful in the healthcare sector where professionals deal with vast amounts of complex information daily, including clinical research papers, patient records, diagnostic reports, and treatment guidelines. The primary goal is to save time and improve efficiency by extracting the most critical information from lengthy texts. Summaries can help doctors, researchers, and other healthcare professionals quickly understand the essence of a study, patient history, or clinical guideline without having to read through the entire document.
TABLE IV: Commonly Used Evaluation Metrics for NLP tasks.
Task Evaluation Metrics
Text Generation BLEU, METEOR, ROUGE, CIDEr, Perplexity
Machine Translation BLEU, TER (Translation Edit Rate), METEOR, chrF++, Adequacy, Fluency
TABLE V: The collection of existing Med-LLMs.
@@ III-A4 Medical Text Generation
Codex-Med [156] 2022 ICL USMLE, MedMCQA,PubMedQA
@@
GPT-4-Med [72] 2023 ICL USML E, MultiMedQA
DeID-GPT [158] 2023 ICL i2b2/UTHealth de-identification task
@@
DoctorGLM [160] 2023 IFT Chinese medical Dialogues
@@
Visual Med-Alpaca [163] 2023 PT, IFT Medical QA
BianQue [164] 2023 IFT Medical QA
Med-PaLM 2 [165] 2023 IFT MultiMedQA, Long-form QA
@@
SoulChat [170] 2023 IFT Long Text, Empathetic Dialogue
@@
ShenNong-TCM-LLM [9] 2023 IFT Chinese Medicine Instruction Dataset
MedicalGPT [7] 2023 PT, RLHF Medical Dialogues
ChatMed [9] 2023 IFT Online Consultation
QiZhenGPT [174] 2023 IFT Pharmaceutical Knowledge&A
Med-ChatGLM [175] 2023 IFT Chinese Medical Knowledge Base
@@
WiNGPT [176] 2024 PT, IFT Medical and General Knowledge C-EVAL, WiNEval
Taiyi-LLM [177] 2024 PT, IFT Biomedical Datasets 13 test sets including BC5CDR
Med-Gemini  [178] 2024 IFT NEJM, USMLE-MM, MMMU-HM, ECG-QA
@@
Zhongjing [180] 2024 PT, SFT, RLHF Medical Books, ChatMed, Medical Wiki CMtMedQA, huatuo-26M

### v2

@@ Abstract
With the advent of Large Language Models (LLMs), medical artificial intelligence (AI) has experienced substantial technological progress and paradigm shifts, highlighting the potential of LLMs to streamline healthcare delivery and improve patient outcomes. Considering this rapid technical progress, in this survey, we trace the recent advances of Medical Large Language Models (Med-LLMs), including the background, key findings, and mainstream techniques, especially for the evolution from general-purpose models to medical-specialized applications. Firstly, we delve into the foundational technology of Med-LLMs, indicating how general models can be progressively adapted and refined for the complicated medical tasks. Secondly, the wide-ranging applications of Med-LLMs are investigated across various healthcare domains, as well as an up-to-date review of existing Med-LLMs. The transformative impact of these models on daily medical practice is evident through their ability to assist clinicians, educators, and patients. Recognizing the importance of responsible innovation, we discuss the challenges associated with ensuring fairness, accountability, privacy, and robustness. Ethical considerations, rigorous evaluation methodologies, and the establishment of regulatory frameworks are crucial for building trustworthiness in the real-world system. We emphasize the need for ongoing scrutiny and development to maintain high standards of safety and reliability. Finally, we anticipate possible future trajectories for Med-LLMs, identifying key avenues for prudent expansion. By consolidating these insights, our review aims to provide professionals and researchers with a thorough understanding of the strengths and limitations of Med-LLMs, fostering a balanced and ethical approach to their integration into the healthcare ecosystem.
@@
Large Language Model, Medical AI, Trustworthiness.
@@ I Introduction
The emergence of foundation models has sparked a transformative wave within the Artificial Intelligence (AI) community in recent years. The success, particularly their remarkable generalization capacity across a myriad of downstream tasks, is attributed to significantly large model sizes and large-scale pre-training on expansive datasets [1, 2]. Such progression has witnessed the extraordinary evolution of large language models (LLMs) within the Natural Language Processing (NLP) society, characterized by a series of seminal advancements, such as BERT [3], T5 [4], and the latest GPT models [5]. Leveraging these powerful models, AI systems can now excel in generating human-like responses, hinting at the potential for highly sophisticated and interactive applications.
@@
Figure 1: Comparison of General LLMs and Medical LLMs. This chart highlights the diversity and specialization for large language models.
@@
§ V, for tree= top_class [Applications, for tree=fill=yellow!45,apply [Medical Diagnosis, Clinical Report Generation, Medical Education, Medical Robotics, Medical Language Translation, apply_wide] ] [Challenges , for tree=fill=yellow!45,apply [Protected Health Information, Clinical Workflows, Safety and Accountability,apply_wide] ] ] [Trustworthiness
@@
Figure 2: Organization of the Survey on Medical Large Language Models. his detailed chart outlines the survey’s structure, covering background, technology, medical tasks and data, evaluation methods, specific medical LLMs, algorithms, applications, trustworthiness and safety, and future directions. It traces the evolution from early NLP to the latest advancements in medical LLMs, including their development, pre-training, fine-tuning, and various evaluation metrics.
@@
Overall, through these advanced explorations in the field of Med-LLMs, various effective perspectives are delivered to promote the rapid development of medical AI societies. From the technical point, it is still a basic manner to improve the performance of Med-LLMs via increased model scale and complexity, which is consistent with the LLM’s scaling law [77]. The Med-LLMs, often pre-trained and fine-tuned on massive biomedical literature corpora like PubMed [78], outperform general-purpose language models on domain-specific tasks. To adapt to diverse clinical tasks, fine-tuning Med-LLMs over an array of specialized medical tasks is required. Essentially, to track the potential pathways for developing Med-LLMs, several reviews [79, 80] primarily focus on exploring the clinical capabilities and applications of Med-LLMs, including Electronic Health Records (EHRs) [81], health education [82], and diagnostic decision making [83].
@@
Figure 3: Evolution of Medical Large Language Models from 2019 to 2024. This roadmap highlights the progression and diversification of medical large language models over the years. Starting with BioBERT in 2019, the field has seen the emergence of specialized models for medical information analysis, Q&A, text generation, and psychological support. Notable developments include ClinicalBERT, BioMegatron, and PubMedBERT in 2020, followed by a surge in models like ChatDoctor, DoctorGLM, and Visual Med-Alpaca in 2022. The roadmap culminates with advanced models such as Med-Gemini and Health-LLM in 2024, reflecting the ongoing innovation in leveraging AI for healthcare applications.
@@ II Background and Technology
II-A Model Architecture
@@
II-A1 Encoder-Decoder Structure
@@
II-A2 Self-Attention Mechanism
@@ II-A3 Positional Encoding
II-A4 Residual Connections & Layer Normalization
For each layer, residual connections are employed to mitigate the problem of vanishing gradients during backpropagation. After the residual connection, layer normalization is applied to stabilize and speed up the training process. This structure ensures that the model can effectively learn from very deep architectures without suffering from degradation issues.
II-B Training Techniques
II-B1 Discussion on Pre-training Stage
@@
$\bullet$ Next Word Prediction (NWP) [87]. Known as Language Modeling, NWP aims to predict the most possible next word in a sequence given the previous words. At its core, NWP relies on understanding the context provided by the preceding words. This context helps the model infer the statistical likelihood of different words following in the sequence, i.e., given a huge corpus, the entire LLM is optimized with maximum likelihood estimation (MLE).
$\bullet$ Mask Language Modeling (MLM) [3]. MLM involves training a model to predict the original identities of certain words (tokens) in a sentence that have been randomly replaced with a special mask token (often “[MASK]”). By predicting the missing words based on remained context, the model can learn to understand the semantic and bidirectional relationships between words. The basic masking strategy involves masking 15% of the tokens in a sequence. Not all masked tokens are treated equally. For the masked tokens, 80% are replaced by the “[MASK]” token, 10% are replaced by a random word, and 10% are left unchanged. This variation prevents the model from simply memorizing the masked positions and forces it to learn more generalizable context representations.
@@
$\bullet$ Replaced Token Detection (RTD) [89]. RTD is a pre-training methodology designed to enhance language models’ capability to discern subtle semantic variations in the text by detecting tokens that have been deliberately substituted with synonyms or semantically equivalent terms. This approach leverages the richness of linguistic synonyms and near-synonyms to create a learning environment that sharpens the model’s sensitivity to the nuances of meaning. During training, two types of sentence instances are generated: (1) Unaltered Sentences: A portion of sentences remain unmodified, serving as a baseline for the model to learn the standard usage of words in context. (2) Altered Sentences: The remaining sentences undergo a careful replacement of selected words with synonyms, creating alternative versions that maintain the overall message but differ in lexical choice. By integrating RTD into the pre-training regimen, models can acquire a more nuanced understanding of semantic similarity and distinction. This refined understanding proves beneficial for downstream tasks that demand high levels of semantic precision, such as text entailment, sentiment analysis, and semantic textual similarity tasks, where the ability to differentiate between subtle shifts in meaning is crucial.
@@
$\bullet$ Next Sentence Prediction (NSP) [3]. Punctuation naturally segments text, offering a logical foundation for developing pre-training strategies based on this inherent structure. The NSP task is designed to train models to discern whether a pair of sentences presented to them are contiguous segments taken from a text corpus. In practice, during NSP’s pre-training phase, two kinds of pairs are constructed: (1) Continuous Pairs (IsNext): Approximately 50% of the time, the model is given sentence pairs that are genuinely consecutive in the original text. (2) Random Pairs (NotNext): For the remaining 50%, the second sentence in the pair is randomly selected from elsewhere in the corpus, forming unrelated or discontinuous sequences. This balanced approach ensures that models not only learn to recognize sequential sentence structures but also develop an understanding of contextual continuity versus disjointed pairings.
@@
$\bullet$ Sentence Order Prediction (SOP) [90]. As a pre-training task, SOP can improve a model’s understanding of the coherence and narrative flow in a text by determining the correct sequential order of sentences, which is particularly useful in training models for tasks that require comprehension of long-form text, dialogue understanding, or any scenario where the sequence of events matters. The primary goal of SOP is to teach models to recognize the logical progression of ideas within a text by predicting whether a pair of sentences are arranged in their original, sequential order or not. Thus, two types of sentence pairs are formulated: (1) Sequential Pairs: Half of the time, sentence pairs are kept in their original order as they appear in the text. These form the positive examples, labeled as ‘IsOrder’. (2) Non-Sequential Pairs: The remaining half are constructed by swapping the order of sentences from different parts of the text, creating incoherent or unrelated sequences. These are labeled as ‘NotOrder’, serving as negative examples. The model is fed these sentence pairs and tasked with predicting whether the given order is correct according to the narrative flow. In this way, SOP encourages models to learn the underlying logic and coherence that connects sentences, enabling them to infer the correct sequence of events or thoughts.
@@
II-B2 Discussion on Fine-tuning Stage
$\bullet$ Supervised Fine-tuning (SFT) [109, 110]. SFT follows a structured approach to adapt pre-trained models to specific tasks by further training them on new, labeled datasets. In SFT, the pre-trained LLM is fine-tuned on labeled datasets using supervised learning techniques.
$\bullet$ Instruction Fine-tuning (IFT) [111]. IFT, also known as prompt-based fine-tuning or instruction tuning, is a recent adaptation of the traditional fine-tuning method, particularly popularized in the context of LLMs. It focuses on teaching models to follow instructions or prompts to perform a variety of tasks without explicit task-specific architecture changes or separate training phases for each task. Instead of fine-tuning on a single-task dataset, instruction fine-tuning involves using a diverse set of instructions or prompts that cover multiple tasks within the same training process. Besides, a medical IFT variant is proposed called Instruction Prompt Tuning (IPT) to train Med-PaLM.
$\bullet$ Parameter-Efficient Fine-ting (PEFT) [112]. PEFT refers to techniques designed to adapt large pre-trained models to specific downstream tasks while minimizing the introduction of new parameters or modifying only a fraction of the existing ones. This approach aims to retain most of the pre-training knowledge, reduce computational costs, and alleviate the risk of overfitting, especially when the target task dataset is small. One popular and classical PEFT strategy is LoRA [113], which introduces low-rank matrices that are multiplied with the weight matrices of specific layers during forward and backward passes. This allows the model to learn task-specific adjustments without changing the original weights directly, reducing the number of new parameters needed.
II-B3 Reinforcement Learning from Human Feedback
II-B4 In-Context Learning
. Also known as few-shot learning for LLMs, ICL [116] refers to the capability of generalizing and adapting to new tasks by providing examples within the prompt, without requiring additional fine-tuning on a specific training dataset. This method leverages the model’s pre-existing knowledge and its ability to understand the context to infer what is being asked and generate appropriate responses.
@@ III-A NLP Tasks under Medical Domain
$\bullet$ Medical Information Extraction. Medical Information Extraction (Med-IE) is a fundamental field to extract relevant medical information from unstructured or semi-structured sources such as electronic health records (EHRs), clinical notes, medical research articles, and other healthcare documents. The aim is to transform this textual data into structured formats that can be easily analyzed, shared, and utilized in various applications like decision support systems, population health management, and medical research. The common applications of Med-IE are summarized in Table II.
@@
$\bullet$ Medical Question Answer. The Medical Question Answering (Med-QA) task is a specific application for building AI systems (e.g., LLMs) capable of understanding and answering complex medical questions from patients, healthcare providers, or researchers. The system typically leverages large amounts of structured and unstructured data sources such as electronic health records, medical textbooks, research articles, and online resources to generate accurate and reliable answers. Here, we summary how the Med-QA system works in the Table II. Information retrieval could involve searching through medical literature, patient histories, or other knowledge bases, which aims to identify and extract key pieces of evidence that are pertinent to the question.
$\bullet$ Medical Natural Language Inference. Medical Natural Language Inference (Med-NLI) is to deal with understanding the relationship between two pieces of medical text and determining whether one sentence logically entails, contradicts, or is neutral with respect to another. This task is also known as Recognizing Textual Entailment (RTE) in the broader context of NLP. In Med-NLI, the inputs are typically pairs of sentences, where one is called the premise and the other is the hypothesis. For example:
@@
$\bullet$ Medical Text Generation Medical Text Generation (Med-Gen), specifically focusing on tasks like text summarization, is an application of artificial intelligence where algorithms process and condense lengthy medical documents into concise summaries. This is particularly useful in the healthcare sector where professionals deal with vast amounts of complex information daily, including clinical research papers, patient records, diagnostic reports, and treatment guidelines. The primary goal is to save time and improve efficiency by extracting the most critical information from lengthy texts. Summaries can help doctors, researchers, and other healthcare professionals quickly understand the essence of a study, patient history, or clinical guideline without having to read through the entire document.
TABLE IV: Overview of Medical Large Language Models (Med-LLMs). This table summarizes key Med-LLMs, detailing their release year, training methods, training data, and evaluation datasets used to assess their performance in various medical tasks.
@@
Codex-Med [148] 2022 ICL Programming Code USMLE, MedMCQA,PubMedQA
@@
GPT-4-Med [72] 2023 ICL General Open-source Data USML E, MultiMedQA
DeID-GPT [150] 2023 ICL General Open-source Data i2b2/UTHealth de-identification task
@@
DoctorGLM [152] 2023 IFT Chinese medical Dialogues -
@@
Visual Med-Alpaca [155] 2023 PT, IFT Medical QA -
BianQue [156] 2023 IFT Medical QA MedDialog-CN,MedDG,CHIPMDCFNPC,IMCS-V2
Med-PaLM 2 [157] 2023 IFT MedMCQA, PubMedQA, etc. MultiMedQA, Long-form QA
@@
SoulChat [162] 2023 IFT Long Text, Empathetic Dialogue SoulChatCorpus,SMILECHAT
@@
ShenNong-TCM-LLM [9] 2023 IFT Chinese Medicine Instruction Dataset -
MedicalGPT [7] 2023 PT, RLHF Medical Dialogues -
ChatMed [166] 2023 IFT Online Consultation -
QiZhenGPT [167] 2023 IFT Pharmaceutical Knowledge&A -
Med-ChatGLM [168] 2023 IFT Chinese Medical Knowledge Base -
@@
WiNGPT [169] 2023 PT, IFT Medical and General Knowledge C-EVAL, WiNEval
Taiyi-LLM [170] 2023 PT, IFT Biomedical Datasets 13 test sets including BC5CDR
Zhongjing [171] 2023 PT, SFT, RLHF Medical Books, ChatMed, Medical Wiki CMtMedQA, huatuo-26M
Med-Gemini  [172] 2024 IFT Medical Knowledge, Clinical Cases NEJM, USMLE-MM, MMMU-HM, ECG-QA